# Detecting Measurement-Induced Entanglement Transitions With Unitary Mirror Circuits

### v1

@@ II.1.3 Proof S8c
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},\Sigma}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\min_{\begin{subarray}{c}\{p,\dotsc,p_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ p_{\min}\leq\mu_{i}\leq\Sigma\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ p_{\min}\leq p\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]p^{\alpha}+\min_{\begin{subarray}{c}\{p,\dotsc,p_{m}\}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma-p\\ p\leq\mu_{i}\leq\Sigma-p\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ p_{\min}\leq p\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-p,p,\Sigma-p}}\right)}}\right)}\Big{]}.\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ p_{\min}\leq p\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-p-\mathopen{}\mathclose{{}\left({m-1}}\right)p}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({m-1}}\right)p^{\alpha}}}\right)}\Big{]}.\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}p\\ p_{\min}\leq p\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left[{\Sigma-mp}}\right]^{\alpha}+mp^{\alpha}}}\right)}\Big{]}.\end{split}$
@@
$\begin{gathered}\frac{\partial}{\partial p}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left[{\Sigma-mp}}\right]^{\alpha}+mp^{\alpha}}}\right)}}\right]=m\alpha\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({p^{\alpha-1}-\mathopen{}\mathclose{{}\left[{\Sigma-mp}}\right]^{\alpha-1}}}\right)>0\\ \Rightarrow F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},\Sigma}}\right)=\mathopen{}\mathclose{{}\left({\Sigma-mp_{\min}}}\right)^{\alpha}+m(p_{\min})^{\alpha}.\end{gathered}$
@@ II.2 Upper bound
$S^{(n)}_{\alpha}\mathopen{}\mathclose{{}\left({\rho_{n}}}\right)=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=1}^{2\bar{n}}\mu_{i}^{\alpha}}}\right],$ (S12)
@@
$\begin{gathered}1-\epsilon^{(n)}_{D}=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=1}^{D}\mu_{i}^{\alpha}}}\right],\qquad\epsilon^{(n)}_{D}=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=D+1}^{2^{\bar{n}}}\mu_{i}^{\alpha}}}\right],\\ p_{D}>p_{D+1}\end{gathered}$
@@
$\begin{split}S^{(n)}_{\alpha}&\leq\max_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{F^{+}\mathopen{}\mathclose{{}\left({D,1-\epsilon^{(n)}_{D},p_{D},1-\epsilon^{(n)}_{D}-(D-1)p_{D}}}\right)+F^{+}\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D,\epsilon^{(n)}_{D},0,p_{D}}}\right)}}\right]\\ &=\max_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{D^{1-\alpha}\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D}}\right)^{1-\alpha}\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}}}\right)^{\alpha}}}\right].\\ &=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{D^{1-\alpha}\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D}}\right)^{1-\alpha}\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}}}\right)^{\alpha}}}\right].\end{split}$ (S13)
@@ II.3 Lower bound
$\begin{split}S^{(n)}_{\alpha}&\geq\min_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{F^{-}\mathopen{}\mathclose{{}\left({D,1-\epsilon^{(n)}_{D},p_{D},1-\epsilon^{(n)}_{D}-(D-1)p_{D}}}\right)+F^{-}\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D,\epsilon^{(n)}_{D},0,p_{D}}}\right)}}\right]\\ &=\min_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({D-1}}\right)(p_{D})^{\alpha}+\lfloor\frac{\epsilon^{(n)}_{D}}{p_{D}}\rfloor(p_{D})^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}-\lfloor\frac{\Sigma}{p_{D}}\rfloor p_{D}}}\right)^{\alpha}}}\right]\\ &\approx\min_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)(p_{D})^{\alpha-1}}}\right].\end{split}$ (S15)
@@
$\begin{split}&\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)(p_{D})^{\alpha-1}}}\right]\geq\\ &\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}-\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}-\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)^{\alpha}}}\right]\mathopen{}\mathclose{{}\left({\frac{p_{D}}{p_{D}^{\max}}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)p_{D}}}\right)(p_{D})^{\alpha-1}}}\right],\end{split}$
and inserting $p_{D}^{\max}=\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)/D$, one finds
$\begin{split}S^{(n)}_{\alpha}&\geq\min_{p_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left[{\frac{D-D^{\alpha}}{1-\alpha}}}\right]^{1-\alpha}\mathopen{}\mathclose{{}\left({\frac{\epsilon^{(n)}_{D}}{\alpha}}}\right)^{\alpha}}}\right].\end{split}$ (S16)

### v2

@@ II.1.3 Proof S8c
$\begin{split}F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},\Sigma}}\right)&=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\min_{\begin{subarray}{c}\{\mu_{1},\dotsc,\mu_{m+1}\}\\ \sum_{i=1}^{m+1}\mu_{i}=\Sigma\\ p_{\min}\leq\mu_{i}\leq\Sigma\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m+1}\mu_{i}^{\alpha}}}\right]\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ p_{\min}\leq{\bar{\mu}}\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]{\bar{\mu}}^{\alpha}+\min_{\begin{subarray}{c}\{\mu_{1},\dotsc,\mu_{m}\}\\ \sum_{i=1}^{m}\mu_{i}=\Sigma-{\bar{\mu}}\\ {\bar{\mu}}\leq\mu_{i}\leq\Sigma-{\bar{\mu}}\end{subarray}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\sum_{i=1}^{m}\mu_{i}^{\alpha}}}\right]}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ p_{\min}\leq{\bar{\mu}}\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m,\Sigma-{\bar{\mu}},{\bar{\mu}},\Sigma-{\bar{\mu}}}}\right)}}\right)}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ p_{\min}\leq{\bar{\mu}}\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha}+\mathopen{}\mathclose{{}\left({\Sigma-{\bar{\mu}}-\mathopen{}\mathclose{{}\left({m-1}}\right){\bar{\mu}}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({m-1}}\right){\bar{\mu}}^{\alpha}}}\right)}\Big{]}\\ &=\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\smashoperator[]{\min_{\begin{subarray}{c}{\bar{\mu}}\\ p_{\min}\leq{\bar{\mu}}\leq\frac{\Sigma}{m+1}\end{subarray}}^{}}\Big{[}{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left[{\Sigma-m{\bar{\mu}}}}\right]^{\alpha}+m{\bar{\mu}}^{\alpha}}}\right)}\Big{]}.\end{split}$
This time we have separated the smallest term $\bar{\mu}$.
@@
$\begin{gathered}\frac{\partial}{\partial\bar{\mu}}\mathopen{}\mathclose{{}\left[{\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left[{\Sigma-m\bar{\mu}}}\right]^{\alpha}+m{\bar{\mu}}^{\alpha}}}\right)}}\right]=m\alpha\operatorname{sign}\mathopen{}\mathclose{{}\left[{1-\alpha}}\right]\mathopen{}\mathclose{{}\left({{\bar{\mu}}^{\alpha-1}-\mathopen{}\mathclose{{}\left[{\Sigma-m\bar{\mu}}}\right]^{\alpha-1}}}\right)>0\\ \Rightarrow F^{-}_{\alpha}\mathopen{}\mathclose{{}\left({m+1,\Sigma,p_{\min},\Sigma}}\right)=\mathopen{}\mathclose{{}\left({\Sigma-mp_{\min}}}\right)^{\alpha}+m(p_{\min})^{\alpha}.\end{gathered}$
@@ II.2 Upper bound
$S^{(n)}_{\alpha}\mathopen{}\mathclose{{}\left({\rho_{n}}}\right)=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=1}^{2^{\bar{n}}}\mu_{i}^{\alpha}}}\right],$ (S12)
@@
$\begin{gathered}1-\epsilon^{(n)}_{D}={\sum_{i=1}^{D}\mu_{i}^{\alpha}},\qquad\epsilon^{(n)}_{D}={\sum_{i=D+1}^{2^{\bar{n}}}\mu_{i}^{\alpha}},\\ \mu_{i}>\mu_{i+1}\end{gathered}$
@@
$\begin{split}S^{(n)}_{\alpha}&\leq\max_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{F^{+}\mathopen{}\mathclose{{}\left({D,1-\epsilon^{(n)}_{D},\mu_{D},1-\epsilon^{(n)}_{D}-(D-1)\mu_{D}}}\right)+F^{+}\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D,\epsilon^{(n)}_{D},0,\mu_{D}}}\right)}}\right]\\ &=\max_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{D^{1-\alpha}\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D}}\right)^{1-\alpha}\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}}}\right)^{\alpha}}}\right].\\ &=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{D^{1-\alpha}\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D}}\right)^{1-\alpha}\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}}}\right)^{\alpha}}}\right].\end{split}$ (S13)
@@
We can further bound this, as
$\begin{split}S^{(n)}&\leq\log 2+\epsilon^{(n)}_{D}\log\mathopen{}\mathclose{{}\left[{2^{\bar{n}}}}\right]+\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)\log D\\ &=\log 2+\epsilon^{(n)}_{D}\log\mathopen{}\mathclose{{}\left[{2^{\bar{n}}/D}}\right]+\log D\\ &\leq\log 2+\mathopen{}\mathclose{{}\left({1-F}}\right)\log\mathopen{}\mathclose{{}\left[{2^{\bar{n}}/D}}\right]+\log D\\ &=\log 2+\bar{n}\mathopen{}\mathclose{{}\left({1-F}}\right)\log 2+F\log D.\end{split}$ (S15)
Using first $0\leq\epsilon^{(n)}_{D}\leq 1$ and then $D\leq 2^{\bar{n}}$ and $\epsilon^{(n)}_{D}\leq 1-F$. With $\bar{n}\leq N/2$ we find the bound used in the main text.
@@ II.3 Lower bound
$\begin{split}S^{(n)}_{\alpha}&\geq\min_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{F^{-}\mathopen{}\mathclose{{}\left({D,1-\epsilon^{(n)}_{D},\mu_{D},1-\epsilon^{(n)}_{D}-(D-1)\mu_{D}}}\right)+F^{-}\mathopen{}\mathclose{{}\left({2^{\bar{n}}-D,\epsilon^{(n)}_{D},0,\mu_{D}}}\right)}}\right]\\ &=\min_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({D-1}}\right)(\mu_{D})^{\alpha}+\lfloor\frac{\epsilon^{(n)}_{D}}{\mu_{D}}\rfloor(\mu_{D})^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}-\lfloor\frac{\Sigma}{\mu_{D}}\rfloor\mu_{D}}}\right)^{\alpha}}}\right]\\ &\approx\min_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)(\mu_{D})^{\alpha-1}}}\right].\end{split}$ (S16)
@@
$\begin{split}&\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)(\mu_{D})^{\alpha-1}}}\right]\geq\\ &\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}-\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}-\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}-\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)^{\alpha}}}\right]\mathopen{}\mathclose{{}\left({\frac{\mu_{D}}{\mu_{D}^{\max}}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left({\epsilon^{(n)}_{D}+\mathopen{}\mathclose{{}\left({D-1}}\right)\mu_{D}}}\right)(\mu_{D})^{\alpha-1}}}\right],\end{split}$
and inserting $\mu_{D}^{\max}=\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)/D$, one finds
$\begin{split}S^{(n)}_{\alpha}&\geq\min_{\mu_{D}}\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left({1-\epsilon^{(n)}_{D}}}\right)^{\alpha}+\mathopen{}\mathclose{{}\left[{\frac{D-D^{\alpha}}{1-\alpha}}}\right]^{1-\alpha}\mathopen{}\mathclose{{}\left({\frac{\epsilon^{(n)}_{D}}{\alpha}}}\right)^{\alpha}}}\right].\end{split}$ (S17)